# Enhancement of quantum sensing in a dissipatively coupled two-mode system

### v1

@@ IV.2 Symmetric broken region
The first-order pole ($m=1$) in the Laurent expansion [Eq. (61)] directly dictates the scaling of the QCRB. Substituting Eq. (61) into the QCRB expression [Eq. (56)], the frequency resolution scales as $\delta\omega\sim(\omega-\omega_{0})$, indicating a linear divergence in sensitivity as $\omega\rightarrow\omega_{0}$.
@@ IV.3 Exceptional Points
V CONCLUSION
@@
In summary, we analyze the sensing sensitivity in a dissipatively coupled two-mode system, which exhibits the anti-PT symmetry under specific conditions. Our investigation covers the anti-PT symmetric unbroken region, the broken region, as well as the EPs and their vicinity. We find that in the unbroken region, only the undetuned critical frequency is present, and the QCRB varies linearly with the probing frequency. In the broken region, the detuned frequency emerges. When the probing frequency is near the detuned frequency, the QCRB decreases nonlinearly and has a valley value, enabling extremely high sensitivity within a narrow range. However, if the probing frequency is not sufficiently close to the detuned frequency, the behavior resembles that of the broken region. In contrast, when the system parameters meet the condition of EPs, the QCRB exhibits a quadratic dependence on the deviation of probing frequency, leading to superior and more robust sensing performance. Our work provides a theoretical foundation for noise-resilient sensing based on anti-PT symmetric systems and avoids the use of post-selection, opening avenues for the design of next-generation quantum sensors.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants No. 12204424, No. 12147149, No. 12565001, No. 12205054), the China Postdoctoral Science Foundation (Grant No. 2022M722889), and the Natural Science Foundation of Jiangxi Province (No. 20252BAC200163).

### v2

@@ IV.2 Symmetric broken region
The first-order pole ($m=1$) in the Laurent expansion Eq. (53) directly dictates the scaling of the QCRB. Substituting Eq. (53) into the QCRB expression Eq. (48), the frequency resolution scales as $\delta\omega\sim(\omega-\omega_{0})$, indicating a linear divergence in sensitivity as $\omega\rightarrow\omega_{0}$.
@@ IV.3 Exceptional Points
V Model Feasibility Analysis
We construct an anti-PT symmetric system by introducing a dissipative coupling channel that interacts individually with two resonators, thereby establishing dissipative coupling between them. This channel, commonly referred to as a memory or bus in physical implementations [77], has been widely adopted in various theoretical models and experimentally validated. Notably, such dissipative coupling is realizable not only in optical cavity systems [78, 79] but also in multimode platforms [55] including optomechanical, superconducting qubit, and cavity magnonic systems.
Figure 6: Dissipative coupling realized in multiple systems. The system schematics: superconducting qubit [80], cavity magnon [64], photonic crystal [81], optomechanics [82], atoms [83], whispering-gallery mode [84].
In studies of optomechanical systems [85, 86], a high-finesse optical cavity formed by a silicon nitride membrane serves as the reservoir. Experimentally, this square membrane ($1\penalty 10000\ \mathrm{mm}\times 1\penalty 10000\ \mathrm{mm}\times 50\penalty 10000\ \mathrm{nm}$) supports closed propagation paths. The reservoir effectively isolates coherent coupling between two optomechanical modes while providing resonant frequencies $\omega_{1}/(2\pi)=788.024\penalty 10000\ \mathrm{kHz}$ and $\omega_{2}/(2\pi)=788.487\penalty 10000\ \mathrm{kHz}$ for the two cavities. Without laser driving, the cavities remain nearly uncoupled and exhibit ultra-low decay rates $\gamma_{1/2}/(2\pi)\sim 1\penalty 10000\ \mathrm{Hz}$. Through precise control of detuning and driving power at $T=4.2\penalty 10000\ \mathrm{K}$, this configuration enables enhanced energy transfer at exceptional points.
For superconducting qubit systems, circuit quantum electrodynamics experiments have demonstrated quantum state transfer from qubits to photons [87, 88]. To achieve scalable architectures, early proposals introduce quantum busses for interqubit connectivity. In a two-qubit implementation, a microwave cavity operating at $\omega_{r}/(2\pi)=5.19\penalty 10000\ \mathrm{GHz}$ functions as a half-wave resonator ($L=\lambda/2=12.3\penalty 10000\ \mathrm{mm}$). When the transition frequency detuning between qubits significantly exceeds the coupling strength ($|\omega_{1}-\omega_{2}|\gg J$), their effective coupling is suppressed, enabling coherent state transfer. This long-range, low-loss interaction provides critical support for on-chip quantum information processing [80]. Multi-qubit extensions utilize transmission lines [89], where individual qubits couple capacitively to the line, with excitation dynamics controlled through spatial arrangement. In electromechanical systems [90], inductive coupling connects two circuits via a transmission line. The construction of bright and dark modes proves essential for satisfying the critical condition $\Gamma_{\mathrm{eff}}\gg k$.
Dissipative coupling through transmission lines has also been implemented in cavity magnonic systems [91, 64, 92]. These configurations typically employ yttrium iron garnet spheres coupled to traveling waveguide modes under magnetic bias. Experimental realizations [65] operate at waveguide frequencies $\omega_{r}\sim\penalty 10000\ \mathrm{GHz}$. When the intrinsic losses of magnons and cavity photons satisfy $\kappa,\gamma\ll g$, the system’s eigenvalues approach zero, ensuring information transfer predominantly through the waveguide. Beyond two-dimensional planar cavities, implementations using three-dimensional and quasi-one-dimensional cavities further confirm experimental feasibility.
The advantages of this model are also evident in experiments coupling optical and acoustic cavities [93]. When a control laser and a signal laser are simultaneously injected into the optical cavity under phase-matching conditions, Brillouin scattering occurs, confining the interaction within a microsphere of approximately $90\penalty 10000\ \mathrm{\mu m}$. The Stimulated Brillouin Scattering coupling between the optical and acoustic cavities is activated by the detuning between the optical cavity’s intrinsic frequency $\omega_{c0}$ and the control laser frequency $\omega_{c}$. The Brillouin scattering spectrum is obtained by scanning the signal laser frequency around $\omega_{s}=\omega_{c}-\Omega_{B}$, a condition corresponding to the parity-time-symmetric state, where $\Omega_{B}$ denotes the acoustic cavity frequency. This approach, benefiting from room-temperature operation and integrability, offers a promising platform for optical storage technologies.
In certain dual-cavity systems, the reservoir also plays a crucial role. When a common reservoir [94] is introduced alongside coherent cavity coupling, directional driving between cavities can be achieved provided the coherent and dissipative coupling strengths satisfy $J=i\Gamma/2$. Conversely, if a strongly dissipative auxiliary mode ($\kappa^{\prime}\gg\kappa$), the resulting dissipation becomes non-directional and reciprocal. Such dissipative coupling can also be realized when both optical cavities simultaneously couple to a waveguide. In the taper-assisted coupling system [95], a dual-bottle resonator and a microsphere resonator are separated by a distance of several micrometers ($\sim 3\penalty 10000\ \mathrm{mm}$). Dissipative coupling also exists between the two control beams of the same chirality, and its strength depends on the atomic diffusion rate [83]. Dissipative coupling can also be realized in whispering-gallery mode resonators [84, 96], either by having two cavities share a common transmission line, or by utilizing the interaction between counter-propagating optical modes in a single cavity. The spatial mode profile within a photonic crystal serves as a key mechanism that induces correlations between subsystems [97].
VI CONCLUSION
@@
In conclusion, we analyze the sensing sensitivity in a dissipatively coupled general two-mode system, which exhibits the anti-PT symmetry under specific conditions. Our investigation covers the anti-PT symmetric unbroken region, the broken region, as well as the EPs and their vicinity. We find that in the unbroken region, only the undetuned critical frequency is present, and the QCRB varies linearly with the probing frequency. In the broken region, the detuned frequency emerges. When the probing frequency is near the detuned frequency, the QCRB decreases nonlinearly and has a valley value, enabling extremely high sensitivity within a narrow range. However, if the probing frequency is not sufficiently close to the detuned frequency, the behavior resembles that of the broken region. In contrast, when the system parameters meet the condition of EPs, the QCRB exhibits a quadratic dependence on the deviation of probing frequency, leading to superior and more robust sensing performance. Our work provides a theoretical foundation for noise-resilient sensing based on anti-PT symmetric systems and avoids the use of post-selection, opening avenues for the design of next-generation quantum sensors.
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grants No. 12204424, No. 12147149, No. 12565001, No. 12205054, No. 12575032, No. 12274376), the China Postdoctoral Science Foundation (Grant No. 2022M722889), the Natural Science Foundation of Henan Province (Grant No. 232300421075), and the Natural Science Foundation of Jiangxi Province (No. 20252BAC200163).
DATA AVAILABILITY
The data that support the findings of this article are openly available [98].